# MatFormer: Nested Transformer for Elastic Inference

### v1

@@ 3.2 Training
MatFormer relies on a simple training strategy of jointly optimizing all the $g$ nested submodels together. To this end, we set the MatFormer loss as a weighted average of loss of $g$ submodels and train for it using the standard stochastic gradient-based optimizers (Shazeer & Stern, 2018):
$\mathcal{L}_{\textsc{joint}}(x,y)=\sum_{i=1}^{g}\lambda_{i}\cdot\mathcal{L}(\mathcal{M}_{i}(x),y),$ (2)
where $\lambda_{i}>0$ is the weight of $i$-th granular submodel. In this paper, we set $\{\lambda_{i}\}_{i=1\dots g}$ to be uniform i.e., $1/g$ but explore tuning $\{\lambda_{i}\}_{i=1\dots g}$ in Appendix D.4 to further improve MatFormer.
The joint training in MatFormer involves one forward pass per each of the $g$ submodels and benefits from portions of shared computation during backpropagation. MatFormer training results in $g$ accurate nested submodels $\mathcal{M}_{1\dots g}$ inside the universal MatFormer model ($\mathcal{M}_{g}$). Note that this simple strategy outperforms various other training techniques (Appendix D.2). Finally, instead of pretraining models with MatFomer structure, we can also induce this structure via finetuning.
MatFormer training is $\sim 15\%$ faster (for $g=4$) than training all the Transformer based equivalent submodels independently (Appendix B). However, MatFormer also enables the extraction of hundreds of smaller submodels along the accuracy-vs-compute curve traced by the $g$ explicitly optimized submodels (Section 3.3). These models emerge for free using Mix’n’Match during inference and drastically reduce the amortized training cost per model obtained through MatFormer. The joint optimization, even without self-distillation from $\mathcal{M}_{g}$, results in smaller submodels that have highly consistent behavior (Section 3.4) with the universal model. Finally, in Appendix B.1, we argue that the training efficiency of MatFormer can be significantly improved through various optimizations.
@@ 3.3 Mix’n’Match
At inference time, it is trivial to extract one of the $g$ submodels $\mathcal{M}_{1}\subset\mathcal{M}_{2}\dots,\subset\mathcal{M}_{g}$ by stacking the corresponding Transformer block $T_{i}$ across layers. However, by selecting different granularity for each MatFormer layer, it is possible to generate a combinatorially large number of accurate smaller models for free. We call this simple procedure Mix’n’Match and observe that these additional model granularities –which were never explicitly optimized – are highly performant.
In fact, we can further increase the number of extracted models by generating interpolating blocks between fixed granulaties (Kusupati et al., 2022). For example, we can generate a $\widetilde{T}$ block that uses first $\frac{1}{2}(m_{i}+m_{i+1})$ neurons in the FFN layer which still tends to be highly accurate.
To summarize, given a computational budget, we can extract a highly accurate model with Mix’n’Match for the constraints rather than using a smaller less accurate model or training a model for this specific constraint (Sections 4.1.14.2). We note that a compute constraint can be satisfied by various Mix’n’Match models with different accuracies, making identifying the best Mix’n’Match configurations without downstream validation is an exciting direction for future work.
@@ 3.4 Deployment
During deployment, all we need to store is the single universal MatFormer model for different types of elastic inference depending on the constraints. In the case of static workloads, where compute resources are known beforehand and the inputs remain relatively similar in difficulty, one can choose the most accurate static submodel for the constraints using Mix’n’Match. This eliminates the usage of a less accurate preexisting model or training of a new one for the specific constraints.
For dynamic workloads, where the compute resources or the input hardness change on the fly, we can use the universal MatFormer model to dynamically extract the optimal submodel for token-based routing in LLMs akin to MoE (Kudugunta et al., 2021; Li et al., 2022) and elastic encoders in dense retrieval (Section 4.2.2). This works largely because all the extracted submodels have high behavioral consistency with universal MatFormer model (Section 4.1) – minimizing the drift across predictions from various submodels. We measure the consistency between two generative models as the percentage of matching tokens generated by them for the same prefix or using the KL divergence of the smaller model outputs with the larger model outputs – this accounts for potential sampling strategies in decoding. This highly consistent nature of MatFormer results in superior inference time speedups for techniques like speculative decoding (Leviathan et al., 2023) (Section 4.1.1) and can assist in reducing prediction drift between cross platform deployments. We also show that higher model consistency also aids metric-space structure preservation in encoder models (Section 4.2.2).
@@ 4 Experiments
In this section, we empirically evaluate MatFormer across modalities (language and vision), model classes (decoder and encoder), and scales (up to 2.6B parameters). Specifically, we train and analyze MatFormer-based decoder-only Language Models – MatLMs (Section 4.1) – and encoder-only Vision Transformers – MatViT (Section 4.2) models with $g=4$ nested granularities across various model sizes. For a fair comparison, we also independently train the Transformer baseline for the submodel of each granularity across model sizes for the same tasks.
We primarily focus on the elastic deployment of MatFormer-based models (Sections 4.1.14.2) for tasks spanning from one-shot generative evals to adaptive image retrieval. Additionally, we also investigate the reliable scaling behavior (Kaplan et al., 2020) of the MatFormer models (Section 4.1.2).
@@ 4.1 MatLM: MatFormer Language Models
We build MatFormer-based decoder-only Language Models – MatLMs – and contrast them to their vanilla Transformer counterparts (LMs) (Liu et al., 2018). The LMs broadly follow the training pipeline and procedure outlined by Thoppilan et al. (2022). For each MatLM model with a set $d_{\text{model}}$, we jointly optimize for $g=4$ nested granularities represented by FFN ratios of $\{0.5,1,2,4\}$ – i.e., only the hidden representation size of the FFN block changes. We denote these submodels as MatLM – {S, M, L, XL} in increasing order of model size and refer to MatLM-XL as the universal MatLM. For baselines, we train vanilla Transformer models with comparable architectures. That is, for each MatLM, we train 4 separate baseline models with FFN ratios of $\{0.5,1,2,4\}$ for a fixed $d_{\text{model}}$ denoted as Baseline – {S, M, L, XL}. We evaluate these models on validation loss (=== log perplexity) and average accuracy on 26 English tasks similar to (Brown et al., 2020; Du et al., 2022; Anil et al., 2023). Of these 26 tasks, we group 5 tasks that require generating multiple tokens under “GEN” and the remaining tasks that involve choosing an option from the input text under “RANK”. Please see Appendix A for further details on training, evaluation, and the datasets.
@@ 4.1.1 Elastic Inference with MatLM
To showcase elastic inference, we evaluate the 2.6B parameter MatLM models on its ability (a) to provide models spanning the accuracy-vs-compute curve using Mix’n’Match (Section 3.3) and (b) to improve post-hoc inference optimization techniques like Speculative Decoding (Leviathan et al., 2023) to further speed-up accurate auto-regressive generation.
Accurate MatLM submodels for every constraint for free with Mix’n’Match. Leveraging Mix’n’Match, a MatLM can provide accurate models for every compute constraint (between S and XL), not just the explicitly optimized granularities {S, M, L, XL}. We evaluate the impact of Mix’n’Match on the 2.6B parameter MatLM in Figure 2 through validation loss and downstream evals and contrast them to four granularities {S, M, L, XL} of the 2.6B baseline LM (all trained independently). In Figures 1(a)1(b)1(c), we show that all MatLM – {S, M, L, XL} models all perform as well as their corresponding baselines – with marginal improvements and drops across the scale.
In Figure 1(a) we see that Mix’n’Match helps obtain many models on the optimal loss-vs-compute curve at zero cost. Moreover, downstream eval tasks on these Mix’n’Match models also mimic this trend, as shown in Figures 1(c)1(b). In a deployment setting that only has $55\%$ of the required compute resources needed for the MatLM-XL model, it is now possible to have a Mix’n’Match submodel with $<2\%$ accuracy drop on RANK evals. Without elastic deployment due to Mix’n’Match, we would see a $>2.5\%$ accuracy drop due to the use of the MatLM-M model. Note that we highlight only a few of the hundreds of accurate Mix’n’Match models along the curves.
(a) Validation loss
(b) 1-shot RANK Evals
(c) 1-shot GEN Evals
(d) Consistency with the XL model
Figure 2: Validation loss & one-shot downstream evaluation scores for the 2.6B MatLM & baseline models. Mix’n’Match helps generate accurate and more consistent models from MatLM that lie on the performance-vs-compute curve spanned by the explicitly optimized submodels.
MatLM submodels speed up speculative decoding. Speculative decoding leverages an accurate lightweight LM as a draft model to autoregressively generate a few tokens, followed by verifying these drafts with a larger model through parallel decoding on the generated tokens. Based on the extent of accurate generation, the draft model is rolled back and reset to the larger model’s output. This results in considerable inference speed-up for the same accuracy as the large model. We point the reader to the original paper for a more detailed explanation (Leviathan et al., 2023).
A draft model that is significantly more consistent with the larger verifier model
Table 1: Inference time speed-ups over a standard 2.6B model through speculative decoding using a 1.5B (S) draft and 2.6B (XL) verifier model.
@@
would lead to less frequent rollbacks of the draft predictions and therefore lower latency. As seen in Figure 1(d) the MatLM submodels can be up to $8.5\%$ more consistent than the baselines to their corresponding XL model. The significant gap persists even in the KL divergence variant of consistency with the XL model’s outputs (see Figure 6 in Appendix). This improved consistency along with the need for only a single universal model positions MatLM favourably to improve techniques that require draft and verifier models such as speculative decoding.
Table 1 shows the inference time speed-ups from speculative decoding using the S and XL submodels of the 2.6B language model for drafting and verification respectively. Speculative decoding with independently trained baseline LMs results in a speed-up of up to $10\%$ over the standard autoregressive decoding of the 2.6B-XL model. But MatLM-based speculative decoding is up to $6\%$ faster than traditional speculative decoding. This additional speed-up can be primarily attributed to the more consistent nature of MatLM-based drafter and verifier models and is further boosted by the ability to share attention cache across models from MatLM which is infeasible for the baselines (see Appendix B.2). Finally, MatLM further reduces the memory overhead for inference by removing the need to have two models during resource-constrained deployment.
(a) Validation loss for XL-models
(b) Validation loss for all models
(c) 1-shot RANK Evals
(d) 1-shot GEN Evals
Figure 3: We train various decoder-only MatLM models at a range of sizes from 78M to 2.6B parameters and observe the scaling trends of all granularities (S, M, L, XL) for validation loss and 1-shot downstream evaluation scores. We find that the MatLM-XL models across scales mimic the training trends of Baseline-XL models. Interestingly, we also note that that validation loss and downstream evaluations follow the scaling trends of the XL-models across all granularities.
@@ 4.1.2 MatLM Scales as well as Vanilla Transformer LMs
Now that we have established that a 2.6B MatLM model and its submodels are as accurate as the baseline Transformer LMs, we want to examine the scalability of training MatLM models. So, we study the scaling properties (Kaplan et al., 2020; Hoffmann et al., 2022) of MatLMs and compare them to vanilla Transformer baseline LMs trained for the same number of tokens. We train models ranging from 78M to 2.6B parameters on 10B to 160B tokens and plot the validation loss for MatLM – {S, M, L, XL} compared against their baselines in Figure 7.
First, in Figure 2(a), we observe that the training of MatLM-XL models across model sizes scale as reliably as the Baseline-XL LMs for loss vs. number of parameters. However, Figure 2(b) interestingly shows that it is not just the XL models but rather all the nested submodels, irrespective of granularity {S, M, L, XL}, of MatLM and Baseline that follow the same scaling trend.
@@
Baseline 20.917 -0.119 1.868
Matformer 17.516 -0.114 1.845
Therefore, we fit a scaling law according to the number of non-embedding parameters ($N$) and training tokens ($D$) for all possible submodels for both MatLMs and the baselines in Table 2. We observe that the fitted parameters are extremely similar, suggesting that MatLMs scale similarly to vanilla Transformer LMs. In Figures 2(c) & 2(d) we also find that the downstream evals for MatLM are within $0.5\%$ of the baselines, with the smaller submodels even outperforming the baselines at scale. Finally, Figure 6(f) in the Appendix shows that the MatLM submodels are more consistent with their XL model compared to the baseline counterparts across scales.
We note that the scaling law equation does not capture how (1) MatLMs have been optimized for multiple submodels and even have performant submodels that have not been explicitly optimized for (Section 4.1.1), and (2) MatLMs and baselines of the same size have different training FLOPs per step. We leave formulations that capture these subtleties to future work and further discuss this in Appendix C.1. We provide full results split by granularity in Appendix C.
4.2 MatViT: MatFormer Vision Transformers
In this section, we extend MatFormer to Vision Transformer (ViT) (Dosovitskiy et al., 2020) based computer vision encoder models. MatFormer-based ViT – MatViT – enables elastic inference for fundamental tasks like image classification and retrieval. To this end, we train the MatFormer variant of the standard ViT-B/16 and ViT-L/16 models – MatViT-B/16 and MatViT-L/16 that are trained with $g=4$ prechosen nested granularities (FFN ratios of $\{0.5,1,2,4\}$). B/16 models are trained on ImageNet-1K (Russakovsky et al., 2015) with AugReg (Steiner et al., 2021) while L/16 models are pretrained on ImageNet-21K (Deng et al., 2009) followed by finetuning on ImageNet-1K. All models are trained with the training setup and optimal hyperparameters of the standard ViT variants from the Scenic library (Dehghani et al., 2022).

### v2

@@ 3.2 Training
MatFormer relies on a simple training strategy of randomly sampling the $g$ nested submodels across training. To this end, for each step we randomly sample a Matformer granularity $i=1,2...,g$ and train for it using the standard stochastic gradient-based optimizers [55]:
$\mathcal{L}_{\textsc{sampling}}(x,y)=\mathcal{L}(\mathcal{M}_{i}(x),y),$ (2)
where $\mathcal{M}_{i}$ is the parameter set of $i$-th granular submodel, with $\mathcal{M}_{i}$ chosen from a probability distribution $\{p_{1},p_{2}...p_{g}\}$. For most experiments in this paper, we uniformly sample each submodel - in Appendix F.3, we find that tuning this probability distribution can result in stronger submodels.
MatFormer training results in $g$ accurate nested submodels $\mathcal{M}_{1\dots g}$ inside the universal MatFormer model ($\mathcal{M}_{g}$), and also enables the extraction of hundreds of smaller submodels along the accuracy-vs-compute curve traced by the $g$ explicitly optimized submodels (Section 3.3). These models emerge for free using Mix’n’Match during inference and drastically reduce the amortized training cost per model obtained through MatFormer. This method results in smaller submodels that have highly consistent behavior (Section 3.4) with the universal model.
@@ 3.3 Mix’n’Match
At inference time, it is trivial to extract one of the $g$ submodels $\mathcal{M}_{1}\subset\mathcal{M}_{2}\dots,\subset\mathcal{M}_{g}$ by stacking the corresponding Transformer block $T_{i}$ across layers. However, by selecting different granularities for each MatFormer layer, it is possible to generate a combinatorially large number of accurate smaller models for free. We call this simple procedure Mix’n’Match and observe that these additional model granularities – which were never explicitly optimized – are highly performant.
For a given compute or parameter budget, there are multiple possible submodels. A common strategy to select an optimal submodel is Neural Architecture Search (NAS) [48, 68]. This, however, is computationally expensive (Appendix D.2). With, Mix’n’Match, we propose gradually increasing the sub-block size with the "least slope". More concretely, we recommend selecting sub-blocks with minimal granularity changes across layers, ensuring that the size of the $j^{th}$ layer is at least that of the $i^{th}$ layer for $j>i$. To give a concrete example, we find that a submodel that uses granularity $g_{2}$ for half the layers then $g_{3}$ for the rest will likely be better than a submodel that uses $g_{1}$ and $g_{4}$ for a similar model size. Our heuristic is underpinned by the training methodology, where each sampled subnetwork maintains consistent layer granularity across the model. Consequently, the model adapts best to configurations where layer granularities are either uniform or display minimal variation. This intution is also backed by NAS, which predicts balanced configurations over skewed (Appendix D.1). Among these "balanced" configurations, we empirically found the increasing with minimum slope configuration to perform the best. In Section 4.1.1 and Appendix D.1, we show that Mix’n’Match works at least as well as using evolutionary search based NAS methods [48] as used by OFA [9].
To summarize, we find that using Mix’n’Match is a simple, cheap and effective hueristic to select a highly performant submodel for a given compute budget (Sections 4.1.14.2). We provide further details and intuition in Appendix D.1.
@@ 3.4 Deployment
The design of MatFormer is beneficial for both static and dynamic workloads:
Static Workloads
To expand upon the example of Llama-2 models in Section 1, a deployment setup might, say, have the latency budget to support 40B parameter Llama model, but can only host a 34B variant because the next bigger model (70B) has significantly higher latency. Training a 40B parameter from scratch would require $4.8*10^{23}$ FLOPs, when training a 34B model and 70B model already cost $4.08*10^{23}$ and $8.4*10^{23}$ FLOPs respectively. So, one would need to settle for a less accurate model despite the larger latency budget. With Matformer, one could obtain a highly accurate 40B model for 0 additional training FLOPS. More precisely, for static workloads, where compute resources are known beforehand and the inputs remain relatively similar in difficulty, one can choose the most accurate static submodel for the constraints using Mix’n’Match.
Dynamic Workloads
For dynamic workloads, where the compute resources or the input hardness change on the fly, we can use the universal MatFormer model to dynamically extract the optimal submodel for each token or query. This works especially well for MatFormer because all the extracted submodels have high behavioral consistency with universal MatFormer model (Section 4.1) – minimizing the drift across predictions from various submodels. We measure the consistency between two generative models as the percentage of matching tokens generated by them for the same prefix or using the KL divergence of the smaller model outputs with the larger model outputs – this accounts for potential sampling strategies in decoding. This high consistency results in superior inference time speedups for techniques like speculative decoding [39] (Section 4.1.1) and can assist in reducing prediction drift between cross platform deployments. We also show that higher model consistency also aids metric-space structure preservation in encoder models (Section 4.2.2). Moreover, given the nested architecture of MatFormer, model colocation can be more memory efficient.
@@ 4 Experiments
In this section, we empirically evaluate MatFormer across modalities (language in Section 4.1 and vision in Section 4.2) and model classes (decoder and encoder). We demonstrate the elastic deployment of MatFormer-based models (Sections 4.1.14.2) for tasks spanning from one-shot generative evals to adaptive image retrieval. Additionally, we also investigate the reliable scaling behavior [29] of MatFormer models (Section 4.1.2).
@@ 4.1 MatLM: MatFormer Language Models
Experiment Setting: We build MatFormer-based decoder-only Language Models – MatLMs – and contrast them to their vanilla Transformer counterparts (LMs) [41]. For each MatLM model with fixed $d_{\text{model}}$, we optimize for $g=4$ nested granularities represented by FFN ratios of $\{0.5,1,2,4\}$ – i.e., only the hidden representation size of the FFN block changes. We denote these submodels as MatLM – {S, M, L, XL} in increasing order of model size and refer to MatLM-XL as the universal MatLM. For baselines, we train vanilla Transformer models with comparable architectures. That is, for each MatLM, we train 4 separate baseline models with FFN ratios of $\{0.5,1,2,4\}$ denoted as Baseline – {S, M, L, XL}. In addition, we adapt OFA  [9] and DynaBERT [27] to our language modeling setup, and compare those to MatFormer at the same model size. We evaluate these models on validation loss and average accuracy on 25 English tasks [8, 22, 3]. We note that no additional memory and compute is used during training these methods compared to independently trained Baselines. Please see Appendix B for further details on training, baselines, evaluation, and the datasets.
(a) Validation loss
(b) 1-shot Evalulation
(c) Consistency w/ XL model
Figure 2: Validation loss & one-shot downstream evaluation scores for the 850M MatLM & baseline models. Mix’n’Match helps generate accurate and more consistent models from MatLM that lie on the performance-vs-compute curve spanned by the explicitly optimized submodels.
Results compared to baselines: To showcase efficacy of MatFormer over baselines, we evaluate 850M MatLM model with the corresponding baseline counterparts in Figure 2.
Overall, in Figures 2(a) and 2(b) we observe all granularity submodels of MatLM outperform their baseline counterparts. Specifically, we find that DynaBERT exhibits a significant 0.01 log perplexity gap compared to MatFormer on the 850M model. The underlying reason is DynaBERT employs joint optimization of all granularities, which yields to fewer gradient updates and hence suboptimal performance compared to MatFormer. DynaBERT would require more than 15% extra compute to perform as well as MatLM. OFA, similar to MatFormer, maintains a single universal model but employs random subnetwork sampling during its training. This leads to sampling fewer models close to S and XL granularity, resulting in inferior performance in this regime. The performance gap manifests as a bell-shaped loss curve (Figure 2(a)), highlighting OFA shortcomings in handling the trade-off between maintaining universal (XL) model quality and model elasticity. Additionally, OFA’s training necessitates complicated NAS strategy for optimal submodel selection. However, using NAS at scale is costly and erroneous, which we further discuss in Appendix D.2. We refer the reader to Appendix B.4 for a more detailed discussion of MatFormer performance compared to baselines, and advantages and downsides of each method.
@@ 4.1.1 Elastic Inference with MatLM
Accurate MatLM submodels for every constraint for free with Mix’n’Match. Mix’n’Match enables the MatLM to deliver accurate models for any compute constraint between S and XL, beyond the fixed granularities {S, M, L, XL}. We assess the efficacy of Mix’n’Match on the 850M parameter MatLM, comparing validation loss and downstream performance against independently trained baseline models {S, M, L, XL}. Figure 2(a) demonstrates that Mix’n’Match achieves optimal loss-vs-compute trade-offs at no additional cost. Additionally, downstream evaluations in Figure 2(b) reinforce this trend. In deployment scenarios with only 55% of the compute resources for a MatLM-XL model, a Mix’n’Match submodel approximates the XL’s performance with only about a 1% accuracy drop, compared to a 2% drop when using the MatLM-M model. This highlights the efficiency of Mix’n’Match in creating numerous optimal models, as exemplified by selected instances along the performance curves.
We experimented with several heuristics to select the best subnetwork, but consistently observed that gradually using larger granularities in deeper layers worked the best (Section 3.3). We find that this heuristic better than the evolutionary search based techniques [48], used in OFA [9] in Figures 2(a)2(b). We also find that applying NAS to MatFormer provides no benefit over Mix’n’Match in Figure  6. We discuss additional details on Mix’n’Match in Appendix D.1.
MatLM submodels speed up speculative decoding. Speculative decoding leverages an accurate lightweight LM as a draft model to autoregressively generate a few tokens, followed by verifying these drafts with a larger model through parallel decoding on the generated tokens. When the draft is inaccurate, the draft model is rolled back and reset to the larger model’s output. This results in considerable inference speed-up for the same accuracy as the large model. We point the reader to the original paper for a more detailed explanation [39].
Slow down of this algorithm stems from cases where the smaller model’s predictions disagree with the larger model. A draft model that is significantly more consistent with the larger verifier model would lead to less rollbacks of the draft predictions and therefore lower latency. As seen in Figure 2(c), MatLM submodels can be up to $11.5\%$ more consistent than the baselines to their corresponding XL model. The significant gap persists even in the KL divergence variant of consistency with the XL model’s outputs (see Figure 7 in Appendix). This improved consistency along with the need for only a single universal model positions MatLM favorably to improve techniques that require draft and verifier models such as speculative decoding.
Table 2: Inference time speed-ups over a standard 850M model through speculative decoding using a 393M (S) draft and 850M (XL) verifier model.
@@
Table LABEL:tab:spec shows the inference time speed-ups from speculative decoding using the S and XL submodels of the 850M language model for drafting and verification respectively. Speculative decoding with independently trained baseline LMs results in a speed-up of up to $10\%$ over the standard autoregressive decoding of the 850M-XL model. But MatLM-based speculative decoding is up to $6\%$ faster than traditional speculative decoding. This additional speed-up can be primarily attributed to the more consistent nature of MatLM-based drafter and verifier models and is further boosted by the ability to share attention cache across models from MatLM which is infeasible for the baselines (see Appendix C.1). Finally, MatLM further reduces the memory overhead for inference by removing the need to have two models during resource-constrained deployment.
@@ 4.1.2 MatLM Scales as well as Vanilla Transformer LMs
(a) Validation loss: XL-models
(b) Validation loss: all models
(c) 1-shot Evals
Figure 3: We train various decoder-only MatLM models at a range of sizes from 78M to 850M parameters and observe the scaling trends of all granularities (S, M, L, XL) for validation loss and 1-shot downstream evaluation scores. We find that the MatLM-XL models across scales mimic the training trends of Baseline-XL models. Interestingly, we also note that that validation loss and downstream evaluations follow the scaling trends of the XL-models across all granularities.
Now that we have established that a 850M MatLM model and its submodels are at least as accurate as the baseline Transformer LMs, we want to examine the scalability of training MatLM models. So, we study the scaling properties [29, 25] of MatLMs and compare them to vanilla Transformer baseline LMs trained for the same number of tokens. We train models ranging from 78M to 850M parameters on 10B to 80B tokens (per granularity) and plot the validation loss for MatLM – {S, M, L, XL} compared against independently trained baselines in Figure 9.
@@
Baseline 14.08 -0.10 0.89
Matformer 21.60 -0.13 1.33
First, in Figure 3(a), we observe that the training of MatLM-XL models across model sizes scale as reliably as the Baseline-XL LMs for loss vs. number of parameters. Figure 3(b) interestingly shows that all granularities {S, M, L, XL}, of MatLM and Baseline follow the same scaling trend. Therefore, we fit a scaling law according to the number of non-embedding parameters ($N$) and training tokens ($D$) for all possible submodels for both MatLMs and the baselines in Table 3. We observe that the fitted parameters are extremely similar, suggesting that MatLMs scale similarly to vanilla Transformer LMs. In Figure 3(c) we also find that the downstream evals for MatLM $0.3\%$ better than the baselines, with the smaller submodels further outperforming the baselines at scale by upto $1.4\%$. Finally, Figure 9(f) in the Appendix shows that the MatLM submodels are more consistent with their XL model compared to the baseline counterparts across scales.
We note that the scaling laws do not capture how MatLMs have been optimized for multiple submodels and even have performant submodels that have not been explicitly optimized for (Section 4.1.1) We leave formulations that capture these subtleties to future work and further discuss this in Appendix E.1. We provide full results split by granularity in Appendix E.
@@
4.2 MatViT: MatFormer Vision Transformers
We extend MatFormer to Vision Transformer (ViT) [21] based computer vision encoder models. MatFormer-based ViT (MatViT) enables elastic inference for fundamental tasks like image classification and retrieval. We train the MatFormer variant of the standard ViT-B/16 and ViT-L/16 models – MatViT-B/16 and MatViT-L/16 that are trained with $g=4$ nested granularities (FFN ratios of $\{0.5,1,2,4\}$). B/16 models are trained on ImageNet-1K [50] with AugReg [57] while L/16 models are pretrained on ImageNet-21K [18] followed by finetuning on ImageNet-1K. All models use the training setup and optimal hyperparameters of standard ViT variants from the Scenic library [16].